# OSMamba: Omnidirectional Spectral Mamba with Dual-Domain Prior Generator for Exposure Correction

### v1

@@ Abstract
footnotetext: Equal contribution. Corresponding author.
@@ 1 Introduction
In real-world scenarios, images often suffer from under- or over-exposure due to complex and variable lighting conditions, inherent limitations of capturing devices, and potential operator errors [1, 4]. For example, under-exposure occurs when image details are obscured in shadows, significantly reducing contrast and clarity; over-exposure can lead to saturation in highlight areas, resulting in irreversible loss of important details. These issues severely affect image quality and can also impact many applications including but not limited to visual recognition [8, 43, 31] and video processing [30, 6, 58]. Consequently, exposure correction [1, 51, 5] has been a critical problem in computer vision, attracting widespread attention from researchers.
@@
In the past decade, many low-light enhancement methods [7, 52, 53, 40, 59, 60] have been developed. They usually assume that input images are purely under-exposed and employ the Retinex theory [45] to decompose images into illumination and reflectance components for separate processing. However, these methods often falter when faced with multi-exposure conditions. The emergence of deep networks [46, 62] has changed the field of exposure correction by enabling end-to-end learning for directly predicting corrected results from inputs. Notably, recent methods [1, 51, 26, 2, 35] incorporate classical image processing algorithms into deep networks for improved performance.
@@
  • We propose a novel exposure correction method OSMamba, which introduces Mamba with an Omnidirectional Spectral Scanning mechanism to effectively model comprehensive long-range dependencies within amplitude and phase spectra, leading to superior illumination correction and structure recovery. To the best of our knowledge, this is the first attempt to introduce the state space model into the exposure correction task.
@@ 2.1 Exposure Correction
A critical challenge in exposure correction is maintaining consistent outputs across inputs with different exposures. Several works have proposed innovative solutions to this problem [25, 27, 28, 34]. ERL [28] explored the relationship between under- and over-exposure optimization by correlating and constraining the correction procedures within a mini-batch. MMHT [34] proposed a Macro-Micro-Hierarchical Transformer with a contrast constraint loss function, alleviating color distortion and inconsistency in correction results. Recent works [9, 26, 2, 36, 35] have also explored new methodologies for exposure correction. LACT [2] introduced an illumination-aware color transform algorithm, focusing on preserving image details. CoTF [35] combined image-adaptive 3D LUTs for global adjustments with pixel-wise transformations for local refinement.
@@ 2.2 State Space Model
Several works [66, 38, 48, 64] have been devoted to developing new vision backbones based on State Space Models (SSMs) [16, 15]. For example, VMamba [38] proposed Visual State Space (VSS) blocks with Cross-Scan for efficient global modeling. RS-Mamba [64] employed an omnidirectional selective scan module, effectively capturing context across various spatial directions in the tasks of remote sensing dense prediction. ZigMa [24] combined Mamba with a Zig-Zag scanning scheme in diffusion models, improving the scalability for high-resolution datasets. Researchers have also explored integrating SSMs with image restoration tasks [49, 18, 33, 65, 67, 54]. FourierMamba [33] introduced frequency scanning in Fourier domain for image deraining, while MambaIR [18] enhanced vanilla Mamba with local modeling and channel attention for image super-resolution and denoising. Moving beyond these works, we propose OSMamba to establish diverse long-range dependencies in the frequency domain through OS-Scan for better illumination correction and structure recovery. Additionally, we develop a DDPG to guide OSMamba to restore details under extreme exposure conditions.
@@ 3.1 Preliminary
State Space Model. The structured state space sequence models (S4) [16] have driven significant advancements in SSMs [19, 20, 3]. Mamba (S6) [15] further utilizes a discretized version of a linear time-varying system that maps an input sequence $\mathbf{x}\in\mathbb{R}^{L\times C}$ to an output sequence $\mathbf{y}\in\mathbb{R}^{L\times C}$ with the same length through a latent state $\mathbf{h}_{k}\in\mathbb{R}^{C\times N}$. The formulation of S6 can be expressed as:
@@ 3.5 Dual-Domain Prior Generator
SCI [40] CVPR 2022 9.97 0.6681 5.837 0.5190 7.49 0.5786 17.86 0.6401 4.45 0.3629 12.49 0.5051 15.96 0.6646
MSEC [1] CVPR 2021 20.52 0.8129 19.79 0.8156 20.08 0.8210 19.62 0.6512 17.59 0.6560 18.58 0.6536 20.38 0.7800
ENC+SID [25] CVPR 2022 22.59 0.8423 22.36 0.8519 22.45 0.8481 21.36 0.6652 19.38 0.6843 20.37 0.6748 22.66 0.8195
@@ 4.1 Experimental Settings
Implementation details. Our method is implemented using the PyTorch framework [44], with all experiments conducted on NVIDIA RTX 4090 GPUs. The network contains 7.5M parameters. We employ progressive training with image patch sizes [192, 256, 320, 384, 512] and corresponding batch sizes [8, 4, 2, 2, 1]. The learning rate follows a cosine annealing schedule [39], decreasing from $2\times 10^{-4}$ to $1\times 10^{-6}$. We use the Adam optimizer [11] with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.99$. The training steps for LCDP, MSEC, and SICE in Stage 1 (Pretraining) are 60K, 100K, and 80K, respectively, followed by 20K joint finetuning steps in Stage 2 (Finetuning). The channel numbers $C$ and $M$ of feature and prior are set to 36 and 256, respectively. For the diffusion model in DDPG, we use a linear noise schedule with $\alpha_{1}=0.99$, $\alpha_{T}=0.1$, and $T=4$.
@@ 4.3 Ablation Study and Discussion
Table 3: Ablation study on the scanning method used in OS-SSM. Cross-Scan [38] No.1-2 represents scanning from top-left to bottom-right and bottom-left to top-right, while No.3-4 represents the reverse directions. OS-Scan No.1-2 corresponds to row-wise and column-wise scanning, while No.3-4 denote positive and negative diagonal scanning.

### v2

@@ Abstract
Equal Contribution. Corresponding Author.
@@ 1 Introduction
In real-world scenarios, images often suffer from under- or over-exposure due to complex and variable lighting conditions, inherent limitations of capturing devices, and potential operator errors [1, 5, 42]. For example, under-exposure occurs when image details are obscured in shadows, significantly reducing contrast and clarity; over-exposure can lead to saturation in highlight areas, resulting in irreversible loss of important details. These issues severely affect image quality and can also impact many applications including but not limited to visual recognition [12, 48, 34] and video processing [33, 7, 63]. Consequently, exposure correction [1, 56, 6] has been a critical problem in computer vision, attracting widespread attention from researchers.
@@
In the past decade, many low-light enhancement methods [11, 57, 58, 45, 64, 65] have been developed. They usually assume that input images are purely under-exposed and employ the Retinex theory [50] to decompose images into illumination and reflectance components for separate processing. However, these methods falter when faced with multi-exposure conditions. The emergence of deep networks [51, 67, 41, 10, 9, 8] has changed the field of exposure correction by enabling end-to-end learning for directly predicting corrected results from inputs. Notably, recent methods [1, 56, 29, 3, 38] incorporate classical image processing algorithms into deep networks for improved performance.
@@
  • We propose a novel exposure correction method OSMamba, which introduces Mamba with an Omnidirectional Spectral Scanning mechanism to effectively model comprehensive long-range dependencies within amplitude and phase spectra, leading to superior illumination correction and structure recovery.
@@ 2.1 Exposure Correction
A critical challenge in exposure correction is maintaining consistent outputs across inputs with different exposures. Several works have proposed innovative solutions to this problem [28, 30, 31, 37]. ERL [31] explored the relationship between under- and over-exposure optimization by correlating and constraining the correction procedures within a mini-batch. Recent works [13, 29, 3, 39, 38] have also explored new methodologies for exposure correction. LACT [3] introduced an illumination-aware color transform algorithm, focusing on preserving image details. CoTF [38] combined image-adaptive 3D LUTs for global adjustments with pixel-wise transformations for local refinement.
@@ 2.2 State Space Model
Several works [70, 43, 53, 69] have been devoted to developing new vision backbones based on State Space Models (SSMs) [20, 19]. For example, VMamba [43] proposed Visual State Space (VSS) blocks with Cross-Scan for efficient global modeling. RS-Mamba [69] employed an omnidirectional selective scan module, effectively capturing context across various spatial directions in the tasks of remote sensing dense prediction. ZigMa [27] combined Mamba with a Zig-Zag scanning scheme in diffusion models, improving the scalability for high-resolution datasets. Researchers have also explored integrating SSMs with image restoration tasks [54, 22, 36, 72, 71, 59, 15]. FourierMamba [36] introduced frequency scanning in Fourier domain for image deraining, while MambaIR [22] enhanced vanilla Mamba with local modeling and channel attention for image super-resolution and denoising. Moving beyond these works, we propose OSMamba to establish diverse long-range dependencies in the frequency domain through OS-Scan for better illumination correction and structure recovery. Additionally, we develop a DDPG to guide OSMamba to restore details under extreme exposure conditions.
@@ 3.1 Preliminary
State Space Model. The structured state space sequence models (S4) [20] have driven significant advancements in SSMs [23, 4]. Mamba (S6) [19] further utilizes a discretized version of a linear time-varying system that maps an input sequence $\mathbf{x}\in\mathbb{R}^{L\times C}$ to an output sequence $\mathbf{y}\in\mathbb{R}^{L\times C}$ with the same length through a latent state $\mathbf{h}_{k}\in\mathbb{R}^{C\times N}$. The formulation of S6 can be expressed as:
@@ 3.5 Dual-Domain Prior Generator
SCI [45] CVPR 2022 9.97 0.6681 5.83 0.5190 7.49 0.5786 17.86 0.6401 4.45 0.3629 12.49 0.5051 15.96 0.6646
MSEC [1] CVPR 2021 20.52 0.8129 19.79 0.8156 20.08 0.8145 19.62 0.6512 17.59 0.6560 18.58 0.6536 20.38 0.7800
ENC+SID [28] CVPR 2022 22.59 0.8423 22.36 0.8519 22.45 0.8481 21.30 0.6645 19.63 0.6941 20.47 0.6793 22.66 0.8195
@@ 4.1 Experimental Settings
Implementation details. Our method is implemented using the PyTorch framework [49], with all experiments conducted on NVIDIA RTX 4090 GPUs. The network contains 7.5M parameters. We employ progressive training with image patch sizes [192, 256, 320, 384, 512] and corresponding batch sizes [8, 4, 2, 2, 1]. The learning rate follows a cosine annealing schedule [44], decreasing from $2\times 10^{-4}$ to $1\times 10^{-6}$. We use the Adam optimizer [35] with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.99$. The training steps for LCDP, MSEC, and SICE in Stage 1 (Pretraining) are 60K, 100K, and 80K, respectively, followed by 20K joint finetuning steps in Stage 2 (Finetuning). The channel numbers $C$ and $M$ of feature and prior are set to 36 and 256, respectively. For the diffusion model in DDPG, we use a linear noise schedule with $\alpha_{1}=0.9$, $\alpha_{T}=0.01$, and $T=4$.
@@ 4.3 Ablation Study and Discussion
Table 3: Ablation study on the scanning method used in OS-SSM. Cross-Scan [43] No.1-2 represents horizontal and vertical scanning starting from the top-left, while No.3-4 represents the reverse directions. OS-Scan No.1-2 corresponds to row-wise and column-wise scanning, while No.3-4 denote positive and negative diagonal scanning.
@@ 5 Conclusion
Acknowledgments. This work was supported in part by National Natural Science Foundation of China (No. 62372016), Guangdong Provincial Key Laboratory of Ultra High Definition Immersive Media Technology (No. 2024B1212010006) and Shenzhen General Research Project under Grant JCYJ20220531093215035.